Write an equation of the line passing through the given points
(2,-17) and (-5,39)

Answer by stanbon(57282)

Write an equation of the line passing through the given points
(2,-17) and (-5,39)
-------------------------
slope = (39--17)/(-5-2) = 56/-7 = -8
--------------------------
intercept:
Form y = mx + b
-17 = -8*2 + b
-17 = -16 + b
b = -1
-------------------------
Equation:
y = -8x -1
=================
Cheers,
Stan H.